OpenStudy (tal_tal):

Fan and medal!!! Taylor bought 6 tickets to a movie. The price of each ticket was the same. After he bought the tickets, his account balance showed a change of −$68.04. What would have been the change to Taylor's account balance had he bought only 1 ticket to the movie? −$11.34 $11.34 −$11.40 $11.40

OpenStudy (tal_tal):

@isaac4321

OpenStudy (tal_tal):

@Nnesha

OpenStudy (tal_tal):

Which fraction is equivalent to -5/9? -9/5 -9/-5 5/9 5/-9

OpenStudy (anonymous):

i think its -11.34

OpenStudy (anonymous):

and the second one is d i believe

OpenStudy (tal_tal):

@isaac4321 thank you!!!!

OpenStudy (anonymous):

np

OpenStudy (anonymous):

can i get a medal

OpenStudy (tal_tal):

@amistre64

OpenStudy (tal_tal):

Is Isaac right?

OpenStudy (amistre64):

isaac is giving out direct answers , so regardless they are breaking the rules of the site

OpenStudy (tal_tal):

Can you help me in a better way please?

OpenStudy (amistre64):

how can we determine a solution? what are your thoughts?

OpenStudy (tal_tal):

divide -$68.04 by 6?

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

and what are your ideas for the second question asked?

OpenStudy (tal_tal):

well I think it's 5/9 because they are opposites

OpenStudy (amistre64):

well, opposites are not equal to each other perse we want to keep the same value, but make it look different. we can either add zero to it, or multiply it by 1

OpenStudy (amistre64):

5/9 is never equal to -5/9 ... nor is it ever equal to 9/5 regardless of the sign

OpenStudy (tal_tal):

So multiply it by 1?

OpenStudy (amistre64):

-5/9 * -1/-1 should give us an equivalent/equal statement

OpenStudy (amistre64):

a convenient form of multiplying by 1 \[\frac{-1}{-1}=1\]

OpenStudy (amistre64):

\[\frac{-5}{9}\frac{-1}{-1}=\frac{?}{?}\]

OpenStudy (tal_tal):

so d?

OpenStudy (amistre64):

5/-9 yes

OpenStudy (tal_tal):

Thank you so much!

OpenStudy (amistre64):

the 'better way' to study and learn is to get you involved and to get your ideas to work with

OpenStudy (amistre64):

good luck :)

OpenStudy (tal_tal):

You're right thank you!